OpenStudy (anonymous):

Find the range of the quadratic function. f(x) = -x2 + 4x - 8 A) [2, ∞) B) [-4, ∞) C) (-∞, -4] D) (-∞, 2]

OpenStudy (anonymous):

find the vertex using -b/(2a)... this is the x-coordinate of the vertex. then you must determine if the parabola opens upwards or downwards... can you give the vertex?

OpenStudy (anonymous):

yeah i think

OpenStudy (anonymous):

(2,-4)

OpenStudy (anonymous):

yes... that's the vertex... now do you know if the parabola opens upwards or downwards?

OpenStudy (s3a):

It opens downwards because the coefficient of x^2 is negative which means the parabola goes infinitely downward hence tha range is from negative infinity to the highest vertical point which is the -4 in (2,-4) so the answer is C.

OpenStudy (anonymous):

thanks

OpenStudy (s3a):

No problem.

OpenStudy (anonymous):

f(x) = -x^2 + 4x - 8 find the vertex (h,k) h = -b/2a = -4/(-2*1)=2 k = -x^2 + 4x - 8 =-2^2 + 4*2 - 8 = -4 + 8 -8=0 so vertex (2,-4) now lets draw it -x^2 will make it open downward